what is the slope of a line perpendicular to the line whose equation is 5x+3y=8?
Answer by MRperkins(300)

the slope of the line perpendicular to 5x+3y=8 is going to be the opposite reciprocal of the slope of 5x+3y=8
Find the slope of 5x+3y=8 and then flip it upside down and give it the opposite sign (positive or negative).
.
Email me with questions
justin.sheppard.tech@hotmail.com